HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

HYPERFINE-RESOLVED SPECTROSCOPIC MODEL AND HIGH-RESOLUTION LINE LIST FOR VANADIUM MONOXIDE (^{51}V^{16}O)

\underline{C. A. BOWESMAN} , S. N. YURCHENKO, J. TENNYSON, Department of Physics and Astronomy, University College London, Gower Street, London WC1E

6BT, UK

Vanadium monoxide (^{51}V^{16}O) is believed to play an important role in the atmospheres of hot-Jupiters, but high-resolution studies have so far failed to detect it, at least in part because of the inaccuracy of available line lists.^{1} It is likely that the large hyperfine splittings in the spectra of VO, arising from the large nuclear spin I = \frac{7}{2} of the ^{51}V atom, has contributed to the non-detections with the current hyperfineunresolved VOMYT line list^{2}. To rectify this, a fully hyperfine-resolved spectroscopic model has been constructed which includes 15 low-lying electronic states (6 quartets and 9 doublets) of VO with the inclusion of hyperfine couplings based on use of the new, hyperfine-resolved version of the diatomic variational nuclear motion program Duo.^{3}

The new spectroscopic model is refined against empirical MARVEL energies derived from experimental transitions,^{4} and hyperfine couplings are fit for the 3 electronic states for which hyperfine effects have been resolved in lab spectra. The new line list comprises over 4 million energy level and 80 billion transitions and is made available via www.exomol.com.

1de Regt et al., Astron. Astrophys., 165, 157 (2023)

2McKemmish et al., Mon. Not. R. Astron. Soc., 463, 771 (2016)

3Qu, Yurchenko and Tennyson, J. Chem. Theory Comput., 18, 1808 (2022)

4Bowesman et al., J. Quant. Spectrosc. Radiat. Transf., 289, 108295 (2022)

p-number: p111  Submitted on Tue, 13 Jun 2023 20:45:49 +0200